# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

### v1

@@ 2.2.2 Multi-modal Deepfake Datasets
Multi-modal deepfake datasets are becoming indispensable as they provide a more intricate testing environment that closely mirrors actual digital interactions, where deceptive content might include any combination of altered videos, synthesized speech, and manipulated text. The multi-modal deepfake datasets simulate the real-world scenarios of how manipulated content delivers misinformation more convincingly. The development of such datasets aims to refine detection algorithms by offering a broad spectrum of data types for analysis. As illustrated in Figure 1, either or both auditory and visual content are forged in audio-visual deepfake dataset, and text-visual deepfake dataset involves text-image pairs where each modality can be falsified.
Audio-Visual Deepfake Datasets Datasets like FakeAVCeleb [5] and Joint Audio-Video Deepfake [10] enhance the depth of deepfake detection research by including video deepfakes with corresponding synthesized, lip-synced audio tracks or by integrating audio alterations into existing visual datasets. These efforts, alongside others like DefakeAVMiT [42] and Trusted Media Challenge (TMC)[43], contribute substantially to building resources capable of addressing the nuanced challenges posed by audio-visual deepfakes. MMDFD[44] extends beyond audio-visual deepfake datasets by encompassing text modality represented as video subtitles, however, without releasing public access thus hindering its further comprehensive evaluation benchmarking.
@@
In general, these efforts[5, 43, 10, 35, 42, 7, 44, 6] encapsulate a significant step toward developing versatile resources for testing and improving detection methods across the complicate multi-modal deepfakes.
@@ 2.3 Evaluation
Evaluation Metrics In prior work involving only the visual modality, where deepfake detection is cast as a classification task, usually a binary classification task, the widely adopted evaluation metrics include $ACC$ (accuracy), $AUC$ (area under ROC curve), $F1-score$, $AP$(average precision), and $EER$ (equal error rate).
Under multi-modal scenarios, where the fine-grained detection for each modality is in a multi-label manner[12, 6], $mAP$ is also reported in addition to the aforementioned metrics.
In recent text-visual grounding works [6], $mIoU$, $IoU_{50}$, $IoU_{75}$ are used as evaluation metrics for visual grounding, where the bounding box of the manipulated region is predicted. Regarding text grounding, where token-level binary prediction is made to indicate the authenticity of each text token, the classification precision and recall are evaluated.
@@
In-domain evaluation refers to evaluating the detectors on testing data that follows the same distribution as training data. In prior works [21, 24, 26], practitioners conduct in-domain evaluation using non-overlapping testing and training data from the same dataset, or in a narrower sense, of the same manipulation type having identical forgery patterns. Under this setting, the evaluation results reflect the efficacy of detectors under a controlled conditions, providing insight into their performance in an idealized environment.
Cross-domain evaluation is the case that testing data and training data are of different distributions, requiring detectors to possess generalizability in real scenarios. The burgeoning generative techniques highlight the necessity of cross-domain evaluation to prevent detectors from overfitting on the specific manipulation types seen during training. There are three major settings to evaluate the cross-domain generalizability:
@@
  • Cross-manipulation: Training and testing occur on the same dataset but involve different manipulation types, e.g., training on FaceSwap manipulations from FF++ and testing on NeuralTextures. This setting requires detailed forgery type annotations besides real/fake labels.
@@ 2.4 History and Challenges of Deepfake Datasets
The evolution of deepfake generation methods has closely paralleled advancements in generative AI technologies, transitioning from early techniques such as GANs and auto-encoders to the more recent and sophisticated Diffusion Models. This progression has necessitated a corresponding shift in deepfake detection strategies, moving from identifying simple visual artifacts to decoding subtler patterns that mirror genuine content, challenging both automated systems and human observers.
Moreover, the emergence of multi-modal deepfake datasets introduces new challenges, requiring detectors to understand and analyze the interactions across different data types, for example, video, audio and text. This integration reflects the real-world scenarios where deepfakes are used, adding layers of difficulty in detection tasks.
In the mean time, the manipulation techniques within deepfakes are also becoming more refined, targeting only specific faces of an image or select frames within a video sequence. This subtlety demands that deepfake detection systems not only identify the presence of fakes but also pinpoint the exact areas of manipulation with high precision.
These developments across various manipulation methods, modalities, and improved qualities highlight the ongoing need for adaptive detection methodologies in response to evolving deepfake datasets.
@@ 3 Method
In the following subsections, we review works predominantly published in the last three years, encompassing three major areas: single-modal detection methods, multi-modal detection methods, and the works related to trust-worthy deepfake detection. Given the extensive array of methods and the intricate technical details inherent in each category, it is impractical to cover all available techniques. Thus, in this survey, we will focus on highlighting key works within each category to delineate the evolutionary trajectory. Figure 3 illustrates our taxonomy of single- and multi-modal deepfake detection in detail.
@@
Figure 3: The taxonomy of deepfake detection methods.
Addition, tier=level5 [[11][152][155][156][157][35][13], tier=last] ] [Attention, tier=level5 [[13][158][159], tier=last] ] [MLP Mixer Layer, tier=level5 [[12], tier=last] ] [Multi-task Strategy, tier=level5 [[10][42][156][12][160],tier=last] ] [Regularization, tier=level5 [[158][159][42][157], tier=last] ] ] ] [Matching-based Learning, tier=level3 [[161][162], tier=last] ] [Others, tier=level3 [[163][164][165][166], tier=last] ] ] [Text-Visual, tier=level2 [[6][167][168][169], tier=last] ] ] ]
@@ 3.1.1 Passive Detection Methods
Passive deepfake detection employs methods and techniques to identify manipulated videos or images based solely on the available content. This approach operates without active intervention or additional information from the creation process. It focuses exclusively on analyzing the existing content to extract artifacts and inconsistencies, without relying on external inputs or markers embedded during the content’s creation. In the following subsections, we review the works that target GAN/VAE generated deepfakes, covering basic and generalized detection methods. For both basic and generalized approaches, we further categorize previous works based on their distinct characteristics.
@@ 3.1.1.1   Basic Detection for GAN/VAE
Advancements continued with Corvi et al.[50] observing that GAN models produce noticeable artifacts during the process. Those artifacts are visible in the Fourier domain and exhibit unusual regular patterns in their auto-correlation. Chen et al. [51] propose using a single patch in conjunction with widely recognized Spatial Rich Model filters to extract artifacts for deepfake detection. Juet et al. [52] introduce the Global and Local Feature Fusion (GLFF) framework to strategically select informative patches for the detailed extraction of local artifacts. Zhang et al. [53] discover that real and synthetically generated images show significantly different latent Gaussian representations when subjected to an inverse diffusion process using a pre-trained diffusion model. By exploiting these distinctions, they can enhance and accentuate subtle artifacts within generated images, improving the effectiveness of their detection methods. Zhang et al. [54] introduce the Cross Appearance-Edge Learning detector, a tool specifically designed to recognize forgery by capturing both coarse and fine-grained appearance and edge patterns at multiple levels of granularity. This method effectively merges knowledge from both the appearance and edge domains to thoroughly expose a broad spectrum of forgery artifacts. Nguyen et al.[55] introduce a multi-task learning system with three key branches: classification, heatmap regression, and self-consistency regression. The heatmap regression and self-consistency regression branches are specially designed to pinpoint vulnerable areas within an image, areas that are particularly indicative of blending artifacts, thereby improving the framework’s ability to detect forgeries.
Consistencies-based Methods The field expanded into temporal artifact detection as researchers recognize the importance of dynamic inconsistencies in videos. Li et al.[56] utilize Long-term Recurrent Convolutional Networks to detect inconsistencies around blinking in videos, while Qi et al.[57] introduce DeepRhythm, which analyzes heartbeat rhythms to spot video forgeries. Haliassos et al.[58] design a spatio-temporal network to detect semantic inconsistencies in mouth movements within forgery videos. The study in [59] introduces a deepfake detection method called latent pattern detection, which identifies temporal inconsistencies across video frames. Haliassos et al.[170] develop a method that employs self-supervised learning to capture detailed video representations by using audio-visual correlations in genuine videos. The detector is then trained on both real and synthetic data to detect facial forgeries and to predict the learned representations from the initial stage for authentic videos. Pang et al.[60] propose MRE-Net consisted of two main components: Bipartite Group Sampling (BGS) and Multiple Rate Branch (MRB). BGS divides the video into bipartite groups to capture various face motion dynamics, while MRB captures short-term and long-term spatial-temporal inconsistencies, respectively. Wang et al.[61] capture spatio-temporal artifacts in one model based on disentangled learning strategy, dividing the convolutional weights of a 3D ResNet into spatial-based and temporal-based groups and updating the two groups alternatively during training. Xu et al.[62] introduce the Thumbnail Layout (TALL) strategy to enhance the ability to capture inconsistencies across frames in fake videos. Their following work TALL++[63] is built on top of TALL[62] by designing a graph reasoning block and introducing a consistency loss to better capture semantic-level inconsistency to improve generalizability. TI2Net [64] uses identity inconsistencies in video frames to authenticate clips. It encodes identity information from each frame into vectors and calculates the temporal identity embedding by analyzing the differences between these vectors. The temporal identity embedding effectively captures identity inconsistencies, which is crucial to determining the authenticity of the video input. In [65], Chen et al. focus on detecting compressed deepfake videos, which often present overlapping compression and tampering artifacts. In a recent work on video deepfake detection [171], Trevine et al. propose a two-stage training pipeline. The first stage uses self-supervised representation learning to enforce audio-visual correspondence with abundant real face videos, combining contrastive learning with a novel audio-visual complementary masking and feature fusion strategy. The second stage involves supervised downstream classification to train a classifier that takes advantage of the lack of cohesion in deepfake videos, thereby enhancing detection accuracy.
@@ 3.1.1.2   Generalized Detection for GAN/VAE
In the realm of deepfake detection, the capacity of generalized detection plays an important role in enhancing the adaptability of the detection methods against a wide range of manipulations. In this subsection, we delve into various categories of approaches such as input level, model level, and learning level methods. Each category provides their own perspectives and techniques to tackle the challenge of generalized deepfake detection.
Input level: Augmentation-based Methods Augmentation-based methods have emerged as a powerful strategy in deepfake detection by enriching training data with complex pseudo-fake examples. Wang et al. [66] develop an attention-based method focused on identifying key facial regions to lay the groundwork for targeted data enhancement. Specifically, the technique involves adaptively masking the facial areas most susceptible to manipulation, as identified through attention analysis. Building on the theme of using real image characteristics, Zhao et al. [67] introduce the Inconsistency Image Generator (I2G). This tool creates well-annotated training datasets by blending genuine images, thereby avoiding artifacts typically introduced by deep neural networks. Further exploring the diversity of forgery characteristics, Chen et al. [68] propose a method to improve the variety of forgeries the model encounters during training. By employing an adversarial training approach, this method not only increases the model’s exposure to diverse forgeries but also improves its ability to detect nuanced differences, making it more adept at identifying sophisticated fake content.
In further exploration, Shiohara et al. [69] utilize synthetic training data through ”self-blended images” (SBIs), which simulate common forgery indicators like blending seams and statistical discrepancies. Larue et al. [70] develop a data augmentation technique that subtly introduces soft discrepancies into synthetic face images. These slight adjustments help train the model, SeeABLE, to better recognize subtle inconsistencies. Li et al. [172] develop a method to enhance detection accuracy for low-quality face forgeries. They introduce a ”bleach generator” that synthesizes a corrective element that, when blended with the input forgery image, produces a ”bleached” image. This approach, bolstered by their image bleach generation technique and two specific loss constraints, effectively improves the detection results.
The prior methods create augmented forgery images through a pixel-level process. Recent works have proposed to conduct the augmentation in latent representation space [71] or feature space [72]. Yan et al. [71] introduce Latent Space Data Augmentation (LSDA), which innovatively manipulates the latent space to generate a broad spectrum of forgery variations. This approach minimizes the risk of overfitting to specific forgery methods and broadens the model’s ability to generalize across various types of manipulated content. In [72], FreqBlender advances pseudo-fake face generation by focusing on frequency knowledge. The approach is based on the hypothesis that faces consist of semantic, structural, and noise information, with structural information most related forgery traces. The authors employ a Frequency Parsing Network to adaptively segment critical frequency components, and create augmented fake samples by adaptively combining those components.
@@
Input Level: Frequency-space based Methods Recently, researchers have expanded beyond RGB space analysis, developing sophisticated techniques that utilize frequency analysis to enhance deepfake detection. Qian et al.[73] highlight the role of frequency analysis in detecting forgery, introducing the Frequency in Face Forgery Network (F3-Net). This innovative two-stream collaborative learning framework leverages frequency-aware decomposed image components and local frequency statistics to identify forgery patterns effectively. Chen et al.[74] develop a face forgery detection method that examines connections between local image regions. They utilize the Multi-scale Patch Similarity Module (MPSM) to detect complex similarity patterns and the RGB-Frequency Attention Module (RFAM) to integrate multi-modal data. Li et al.[75] create an adaptive frequency feature generation module to dynamically identify forgery patterns in images. Their approach also introduces a single-center loss (SCL) technique to reduce intra-class variations among genuine faces and enhance class distinctions, improving feature discrimination in face forgery detection.
Building on these foundational works, subsequent research has introduced advanced approaches with additional complexities. Luo et al.[76] propose a method that leverages high-frequency knowledge through three modules: multi-scale high-frequency feature extraction, residual-guided spatial attention, and cross-modality attention, aimed at enhancing forgery detection by focusing on manipulation signs and harmonizing data aspects. Liu et al.[77] introduce Spatial-Phase Shallow Learning (SPSL), a method that utilizes up-sampling artifacts common in forged content. SPSL combines spatial analysis with phase spectrum insights, emphasizing local texture information to detect forgeries.
@@
Further, Woo et al.[78] present the Attention-based Deepfake detection Distiller (ADD), which applies frequency domain learning and optimal transport theory in knowledge distillation. ADD uses frequency attention distillation to recover high-frequency components lost during compression and multi-view attention distillation to generate multiple attention vectors. Jeong et al.[79] create frequency-level perturbation maps to blur the distinctions between authentic and fabricated images. Their method, similar to adversarial learning, refines both the deepfake detector and the perturbation generator, enhancing the model’s ability to recognize forgeries from various GAN models, although adding complexity and computational demands. In [80], Wang et al. propose a Spatial-Frequency Dynamic Graph method using dynamic graph learning to analyze relation-aware features across spatial and frequency domains. Tan et al.[81] introduce FreqNet, a deepfake detection network that integrates frequency domain learning with a CNN classifier to significantly improve generalization across different scenarios. By applying convolutional layers to both phase and amplitude spectra, their approach detects a broad range of forgery indicators within the frequency domain. Doloriel et al.[82] find that frequency masking improves the generalization of deepfake detection across various generative models, achieving notable performance improvements.
@@
Model Level: Transformer-Based Approaches Recently, transformers [8] have been employed in both static deepfake detection and temporal deepfake segmentation. Wang et al.[83] provide an overview (as of February 28, 2024) of fourteen works adopting Transformer architecture for deepfake detection.
The approach outlined in Saha et al. [84] utilizes a Transformer-based architecture for segmenting deepfake videos, specifically targeting frames that have been manipulated. This method integrates a Vision Transformer (ViT) to analyze spatial features using a Scaling and Shifting (SSF) strategy, alongside a Timeseries Transformer (TsT) that focuses on temporal aspects. The spatial data processed by the ViT is segmented into overlapping windows, which are then analyzed temporally by the TsT to detect deepfake manipulations effectively. In [85], Bai et al. introduce the Action-Units Relation Learning framework, which utilizes a newly designed Action Units Relation Transformer (ART) to enhance forgery detection by analyzing relationships between facial action units (AUs). ART incorporates both AU-specific and AU-agnostic branches to effectively uncover forgery clues. Furthermore, the framework includes the Tampered AU Prediction (TAP) technique, which manipulates AU-related regions to create challenging training samples, thereby refining the detection process.
@@
Using advances in parameter-efficient tuning methods such as Adaptors[86, 88], Mixture of Experts[89, 88], and LoRA[89, 88], innovative strategies are being proposed to integrate these techniques into large pre-trained Vision-Transformers that serve as foundation models. These strategies elegantly address the data-demanding issue while adapting foundation ViTs to deepfake detection and circumvent distortion of originally learned rich knowledge during pre-training. The study by Nguyen et al. [90] investigates the effectiveness of self-supervised pre-trained vision transformers compared to supervised pre-trained transformers and traditional ConvNets. The findings suggest that transformers trained with self-supervised learning on broad datasets, not specifically tailored for deepfake detection, can generally outperform those trained through supervised methods. The DeepFake-Adapter [86] enhances a ViT by adding lightweight, dual-level adapter modules without altering the core architecture. In this method, Shao et al. propose to incorporate Globally-aware Bottleneck Adapters alongside the MLP layers to identify broad forgery cues and uses Locally-aware Spatial Adapters to interact with ViT features, enhancing the detection of localized anomalies in deepfake content. Similarly, Luo et al.[87] propose a novel Forgery-aware Adaptive Vision Transformer to boost generalization. They adapt ViT to deepfake detection by adding two specially designed components and keeping the vanilla ViT frozen to circumvent overfitting issue. The proposed framework is a dual-branch architecture, the CNN branch with Local-aware Forgery Injector aims to inject local information into ViT branch, and ViT-branch comprises vanilla ViT and Global-aware Forgery Adapter targets to learn generalizable features for deepfake detection. Liu et al.[89] propose the use of a mixture of low-rank experts to fully harness the potential of CLIP-ViT, thereby preserving existing knowledge while simultaneously expanding the model’s capacity for transferable deepfake detection. Kong et al. develop a novel method, MoE-FFD [88], which enhances face forgery detection by integrating LoRA and Adapter with the Vision Transformer. This approach effectively identifies forgery clues on both global and local scales. Additionally, their innovative Mixture-of-Experts modules boost the model’s capacity and adaptability, allowing for improved detection performance in a plug-and-play manner across different transformer architectures.

### v3

@@ 2.2.2 Multi-modal Deepfake Datasets
Multi-modal deepfake datasets are becoming indispensable as they provide a more intricate testing environment that closely mirrors actual digital interactions, where deceptive content might include any combination of altered videos, synthesized speech, and manipulated text. The multi-modal deepfake datasets simulate real-world scenarios of how manipulated content delivers misinformation more convincingly. The development of such datasets aims to refine detection algorithms by offering a broad spectrum of data types for analysis. As illustrated in Figure 2, either or both auditory and visual content is forged in audio-visual deepfake datasets, and text-visual deepfake datasets involve text-image pairs where either modality can be falsified.
Audio-Visual Deepfake Datasets Datasets such as FakeAVCeleb [5] and Joint Audio-Video Deepfake [10] improve the depth of deepfake detection research by including video deepfakes with corresponding synthesized, lip-synced audio tracks or by integrating audio alterations into existing visual datasets. These efforts, along with others such as DefakeAVMiT [45] and Trusted Media Challenge (TMC)[46], contribute substantially to building resources capable of addressing the nuanced challenges posed by audio-visual deepfakes. MMDFD[47] extends beyond audio-visual deepfake datasets by encompassing the text modality represented as video subtitles, however, without releasing public access thus hindering its further comprehensive evaluation benchmarking. PolyGlotFake[36] enriches the language diversity of the audio-visual deepfakes, characterized as the first multi-lingual deepfake dataset covering $7$ languages in acoustic content.
@@
In general, these aforementioned efforts, such as [5, 46, 10, 35, 45, 7, 47, 6], encapsulate a significant step toward developing versatile resources for testing and improving detection methods across the complicate multi-modal deepfakes.
@@ 2.3 Evaluation
Evaluation Metrics In prior work involving only the visual modality, where deepfake detection is usually cast as a binary classification task, the widely adopted evaluation metrics include $ACC$ (accuracy), $AUC$ (area under ROC curve), $F1-score$, $AP$(average precision), and $EER$ (equal error rate). Under multi-modal scenarios, where the fine-grained detection for each modality is in a multi-label manner[12, 6], $mAP$ (mean Average Precision) is also reported in addition to the aforementioned metrics.
In recent text-visual grounding works [6], $mIoU$ (mean Intersection over Union), $IoU_{50}$, $IoU_{75}$ are used as evaluation metrics for visual grounding, where the bounding box of the manipulated region is predicted. Regarding text grounding, where a token-level binary prediction is made to indicate the authenticity of each text token, the classification precision and recall are evaluated.
@@
In-domain evaluation refers to evaluating the detectors on the testing data following the same distribution as the training data. In previous work such as [21, 24, 26], practitioners perform an in-domain evaluation using non-overlapping testing and training data from the same dataset, or in a narrower sense, of the same manipulation type having identical forgery patterns. Under this setting, the evaluation results reflect the efficacy of detectors under a controlled conditions, providing insight into their performance in an idealized environment.
Cross-domain evaluation refers to the scenario where the test data and training data come from different distributions, requiring detectors to have generalizability in real-world situations. The evolving generative techniques highlight the necessity of cross-domain evaluation to prevent detectors from overfitting on the specific manipulation types seen during training. There are three major settings to evaluate the cross-domain generalizability:
@@
  • Cross-manipulation: Training and testing occur on the same dataset but involve different manipulation types, e.g., training on FaceSwap manipulations from FF++ and testing on NeuralTextures also from FF++. This setting requires detailed forgery type annotations besides real/fake labels.
@@ 2.4 History and Challenges of Deepfake Datasets
The evolution of deepfake generation methods has closely paralleled advancements in generative AI technologies, transitioning from early techniques such as GAN/VAE[21, 24, 22, 25] to the more recent and sophisticated DM[32, 33, 34]. This progression has necessitated a corresponding shift in deepfake detection strategies, moving from identifying simple visual artifacts to decoding subtler patterns that mirror genuine content, challenging both automated deepfake detectors and human observers.
Moreover, the emergence of multi-modal deepfake datasets[5, 35, 7, 36, 6] introduces new challenges, requiring detectors to understand interactions across different data types, for example, video, audio, and text. This integration reflects the real-world scenarios where deepfakes are used, adding layers of difficulty in detection tasks.
In the mean time, the manipulation techniques within deepfakes are also becoming more refined, targeting only specific facial regions of an image[29, 30, 6] or selecting frames within a video sequence[35, 7]. This subtlety demands that detection systems not only identify the presence of fakes, but also pinpoint the exact areas of manipulation.
In summary, these developments across various manipulation methods , modalities, and improved qualities highlight the ongoing need for adaptive detection methodologies in response to evolving deepfake datasets.
@@ 3 Method
In the following subsections, we review the works published predominantly in the last three years, encompassing the following major areas: single-modal detection methods, multi-modal detection methods. Given the extensive array of methods and the intricate technical details inherent in each category, it is impractical to cover all available techniques. Thus, in this survey, we will focus on highlighting key works within each category to delineate the evolutionary trajectory. Figure 3 illustrates our taxonomy of single- and multi-modal deepfake detection in detail.
@@
Figure 3: The taxonomy of facial deepfake detection methods.
Addition, tier=level5 [[11][169][173][174][175][35][13], tier=last] ] [Attention, tier=level5 [[13][176][177], tier=last] ] [MLP Mixer Layer, tier=level5 [[12], tier=last] ] [Multi-task Strategy, tier=level5 [[10][45][174][12][178],tier=last] ] [Regularization, tier=level5 [[176][177][45][175], tier=last] ] ] ] [Matching-based Learning, tier=level3 [[179][180][181], tier=last] ] [Others, tier=level3 [[182][183][184][185][186], tier=last] ] ] [Text-Visual, tier=level2 [[6][187][188][189], tier=last] ] ] ]
@@ 3.1.1 Passive Detection Methods
Passive deepfake detection employs methods and techniques to identify manipulated videos or images based solely on the available content. This approach operates without active intervention or additional information from the creation process. It focuses exclusively on analyzing the existing content to extract artifacts and inconsistencies, without relying on external inputs or markers embedded during the content’s creation. In the following subsections, we first review the works that target GAN/VAE generated deepfakes, covering basic and generalized detection methods. For both basic and generalized approaches, we further categorize previous works based on their distinct characteristics. Then, we review the works about detecting DMs-generated deepfakes and sequential deepfake detection.
@@ 3.1.1.1   Basic Detection for GAN/VAE
In deepfake detection, artifacts usually refer to unintended irregularities that appear when creating or manipulating digital media using AI technologies. These imperfections can include mismatches in lighting and color [48], unusual facial textures [49], frequency domains [53], inconsistent edge blending [58], among others, that are not typically present in genuine media. In addition, inconsistencies can manifest as temporal anomalies in video sequences, where facial movements [62], identities [71], or noise patterns [72, 73], among others, do not maintain continuity between frames. Detecting these artifacts and inconsistencies is crucial to identifying deepfakes, as they serve as important indicators that the input media has been altered or synthetically generated.
@@
Advancements continued with Corvi et al.[53] observing that GAN models produce noticeable artifacts during the process. Those artifacts are visible in the Fourier domain and exhibit unusual regular patterns in their auto-correlation. Chen et al. [54] propose using a single patch in conjunction with widely recognized Spatial Rich Model filters to extract artifacts for deepfake detection. Juet et al. [55] introduce the Global and Local Feature Fusion (GLFF) framework to strategically select informative patches for the detailed extraction of local artifacts. Zhang et al. [56] discover that real and synthetically generated images show significantly different latent Gaussian representations when subjected to an inverse diffusion process using a pre-trained diffusion model. Using these distinctions, they can enhance and accentuate subtle artifacts within generated images, improving the effectiveness of their detection methods. Zhang et al. [57] introduce the Cross Appearance-Edge Learning detector, a tool specifically designed to recognize forgery by capturing both coarse- and fine-grained appearance and edge patterns at multiple levels of granularity. This method effectively merges knowledge from both the appearance and edge domains to thoroughly expose a broad spectrum of forgery artifacts. Nguyen et al.[58] introduce a multi-task learning system with three key branches: classification, heatmap regression, and self-consistency regression. The heatmap regression and self-consistency regression branches are specially designed to pinpoint vulnerable areas within an image, areas that are particularly indicative of blending artifacts, thereby improving the framework’s ability to detect forgeries.
Consistencies-based Methods The field expanded into temporal artifact detection as researchers recognize the importance of dynamic inconsistencies in videos. Li et al.[59] utilize Long-term Recurrent Convolutional Networks to detect inconsistencies around blinking in videos, while Qi et al.[60] introduce DeepRhythm, which analyzes heartbeat rhythms to spot video forgeries. In [61], Zheng et al. propose to leverage temporal coherence for video face forgery detection. Their framework includes two stages: a fully temporal convolution network that enhances temporal feature extraction by reducing the spatial convolution kernel size, and a Temporal Transformer network that explores long-term temporal coherence. Haliassos et al.[62] design a spatio-temporal network to detect semantic inconsistencies in mouth movements within forgery videos.
To further address temporal inconsistencies, various methods have been developed and applied to domains beyond pixel analysis. Choi et al. [63] propose detecting fake videos by analyzing the inconsistent temporal behavior of style latent vectors between real and fake samples. They observe that generated facial videos exhibit distinct temporal changes in these vectors during stable video generation. In their method, Choi et al. employ a StyleGRU module, trained with contrastive learning, to capture the dynamic properties of style latent vectors. The study in [64] introduces latent pattern detection, which identifies temporal inconsistencies across video frames. Haliassos et al.[65] develop a method that employs self-supervised learning to capture detailed video representations by using audio-visual correlations in genuine videos. The detector is then trained on both real and synthetic data to detect facial forgeries and to predict the representations from the initial stage for authentic videos.
Expanding on the identification of spatio-temporal inconsistencies, Pang et al.[66] propose MRE-Net consisted of two main components: Bipartite Group Sampling (BGS) and Multiple Rate Branch (MRB). BGS divides the video into bipartite groups to capture various face motion dynamics, while MRB captures short-term and long-term spatial-temporal inconsistencies, respectively. Wang et al.[67] capture spatio-temporal artifacts in one model based on a disentangled learning strategy, dividing the convolutional weights of a 3D ResNet into spatial-based and temporal-based groups and updating the two groups alternatively during training. Based on the observation that fake videos struggle to maintain natural spatio-temporal consistency under various forgery methods, Zhang et al. [68] introduce the Natural Consistency method for detecting deepfake videos. Their approach leverages CNNs to extract spatial features and Transformers to learn long-range spatiotemporal representations. Additionally, they introduce a Spatial Predictive Module to predict random masked spatial features and a Temporal Contrastive Module to regularize latent distances by shuffling the natural order, both of which effectively enhance detection accuracy.
Several methods focus on using other inconsistencies, such as semantic, identity, and noise patterns, to enhance detection capabilities. Xu et al.[69] introduce the Thumbnail Layout (TALL) strategy to enhance the ability to capture inconsistencies between frames in fake videos. Their following work TALL++[70] is built on top of TALL[69] by designing a graph reasoning block and introducing a consistency loss to better capture semantic-level inconsistency to improve generalizability. TI2Net [71] uses identity inconsistencies in video frames to authenticate clips. It encodes identity information from each frame into vectors and calculates the temporal identity embedding by analyzing the differences between these vectors. The temporal identity embedding effectively captures identity inconsistencies, which is crucial to determining the authenticity of the video input. Based on the assumption that background squares remain unmodified in fake samples, Wang et al. [72] propose utilizing the consistency of noise patterns between the face regions and background patches to differentiate real and fake frames. In [73], Chen et al. focus on detecting compressed deepfake videos, which often present overlapping compression and tampering artifacts.
@@ 3.1.1.2   Generalized Detection for GAN/VAE
In the realm of deepfake detection, the generalizability of detection plays an important role in enhancing the adaptability of the detection methods against a wide range of manipulations. In this subsection, we delve into various categories of approaches, such as input level, model level, and learning level methods. Each category provides its own perspectives and techniques to tackle the challenge of generalized deepfake detection.
Input level: Augmentation-based Methods Augmentation-based methods enhance deepfake detection by enriching training datasets with complex pseudo-fake examples. Wang et al. [74] develop an attention-based method that targets key facial regions for data enhancement by adaptively masking areas most likely to be manipulated. Zhao et al. [75] introduce the Inconsistency Image Generator (I2G), creating well-annotated datasets by blending genuine images to avoid common artifacts. Chen et al. [76] further diversify forgery characteristics in training, using adversarial methods to expose the model to a wider array of forgeries and enhance detection capabilities. In further exploration, Shiohara et al. [77] employ “self-blended images” (SBIs) to simulate typical forgery indicators like blending seams. Larue et al. [78] introduce soft discrepancies into synthetic face images, aiding the model in recognizing subtle inconsistencies. Li et al. [190] improve low-quality forgery detection with a “bleach generator” that corrects and enhances input forgeries for better accuracy.
Traditional methods typically involve pixel-level augmentations. However, recent approaches [79, 80] explore augmentations in latent and frequency spaces. Yan et al. [79] use Latent Space Data Augmentation (LSDA) to generate diverse forgery variations, minimizing overfitting and enhancing generalization. FreqBlender [80] focuses on frequency knowledge, employing a Frequency Parsing Network to segment and combine critical frequency components for more authentic pseudo-fakes.
@@
Input Level: Frequency-space based Methods Recently, researchers have expanded beyond RGB space analysis, developing sophisticated techniques that utilize frequency analysis to enhance deepfake detection. Qian et al.[82] introduce the Frequency in Face Forgery Network (F3-Net), utilizing frequency-aware components and statistics to identify forgeries. Chen et al.[83] analyze local region connections using the Multi-scale Patch Similarity Module (MPSM) and RGB-Frequency Attention Module (RFAM) to detect complex patterns and integrate multi-modal data. Li et al.[84] develop an adaptive frequency feature generation module, incorporating a single-center loss (SCL) to distinguish class features effectively.
@@
Building upon these foundational works, subsequent research has introduced advanced approaches with additional complexities. Luo et al.[85] enhance forgery detection with a focus on high-frequency knowledge and cross-modality attention, optimizing data harmonization. Liu et al.[86] present Spatial-Phase Shallow Learning (SPSL) that combines spatial and phase spectrum analysis to spotlight up-sampling artifacts in forged content. Further, Woo et al.[87] present the Attention-based Deepfake detection Distiller (ADD), which applies frequency domain learning and optimal transport theory in knowledge distillation. ADD uses frequency attention distillation to recover high-frequency components lost during compression and multi-view attention distillation to generate multiple attention vectors. Jeong et al.[88] create frequency-level perturbation maps that blur distinctions between genuine and fake images, refining both the detector and perturbation generator. Wang et al.[89] propose the Spatial-Frequency Dynamic Graph method using dynamic graph learning to analyze relation-aware features. Tan et al.[90] develop FreqNet, integrating frequency domain learning with CNN classifiers to enhance detection across various scenarios. Doloriel et al.[91] demonstrate how frequency masking can significantly improve the generalization of deepfake detection, achieving notable performance enhancements.
@@
Model Level: Transformer-Based Approaches Recently, transformers [8] have been employed in both static deepfake detection and temporal deepfake segmentation. Wang et al.[92] provide an overview (as of February 28, 2024) of fourteen works adopting the Transformer architecture for deepfake detection. Building on these advancements, Saha et al. [93] utilize a Transformer-based architecture, integrating a Vision Transformer (ViT) with a Timeseries Transformer (TsT) to focus on manipulated frames, enhancing detection capabilities by analyzing spatial and temporal features. Expanding on the use of transformers, Bai et al. [94] introduce the Action Units Relation Transformer within their Action-Units Relation Learning framework to analyze relationships between facial action units, enhancing forgery detection through AU-specific and agnostic branches.
Further advancing in this direction, Zhang et al.’s UMMAFormer [95] employs the Temporal Feature Abnormal Attention module and a Parallel Cross-Attention Feature Pyramid Network for refined forgery segment prediction. Pellicer et al. [96] propose the Prototype-based Unified Framework, using features from a pre-trained DINOV2 model [191] for prototype learning, aiming to improve deepfake detection through a Prototype Learning layer that clusters and compares prototypes.
@@
Using advances in parameter-efficient tuning methods such as Adaptors[97, 99], Mixture of Experts[100, 99], and LoRA[100, 99], innovative strategies are being proposed to integrate these techniques into large pre-trained Vision-Transformers that serve as foundation models. These strategies elegantly address the data-demanding issue while adapting foundation ViTs to deepfake detection and circumvent distortion of originally learned rich knowledge during pre-training. These techniques efficiently address data demands and prevent distortion of pre-trained knowledge [101]. The DeepFake-Adapter [97] introduces dual-level adapters to ViTs, enhancing anomaly detection without core architecture modifications. Luo et al. [98] develop a Forgery-Aware Adaptive Vision Transformer, employing a dual-branch architecture to boost feature generalization and prevent overfitting. Liu et al.[100] utilize a mixture of low-rank experts to maximize the potential of CLIP-ViT for deepfake detection, aiming to preserve pre-trained knowledge while enhancing the model’s adaptability. Similarly, Kong et al. introduce MoE-FFD [99], a method that incorporates LoRA and Adapter with the Vision Transformer to improve the detection of face forgeries.